FIND THE P VALUE

Consider the following hypotheses:

 

H0: p ≥ 0.50

HA: p < 0.50

 

Compute the p-value based on the following sample information. (You may find it useful to reference the appropriate table: z table ort table) (Round "z" value to 2 decimal places. Round intermediate calculations to at least 4 decimal places and final answers to 4 decimal places.)


a.x = 40; n = 87

b. x = 108; n = 269
Answer #1

solution Here we from p-value. have to find the given data: Ho: P >0.50 Hai P<0.50 a) Given that x=40 and n=87 P = k/n = 40/2b) Given that x=108 and n =269 È = */n : 108/269 -10.40 test statistic za - n 0.40 - Oiso [ 0.50 (-0.50) V 269 50.0009 ? -0.12: 0:47-0.50 0.50 (1-0.50) 53 -0.03 v0.0047 -0.03 0.0686 : F0.4373] 1 -10.44 - P(Z <-0.44)= 10.3300 -d) Given that P 2047 and